OpenStudy (anonymous):

Simplify the expression. 22+(3^2 -4^2)

OpenStudy (anonymous):

22+(9-16) Then simplify this.

OpenStudy (anonymous):

huh

OpenStudy (cwrw238):

compute whats in the parentheses first

OpenStudy (cwrw238):

what is 9 - 16?

OpenStudy (anonymous):

PEMDAS 22 + (3^2 - 4^2) Parenthesis. Exponents. (3^2 - 4^2) 3^2 = 3 * 3 = 9 4^2 = 4 * 4 = 16 9 - 16 = 22 + (9 - 16)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome. (: